### v1

# Varying Newton constant,entropy and the black hole evaporation law

Abstract
In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of an ideal fluid plus the cosmological term. We consider time-dependent Newton “constant” $G$, the cosmological term $\Lambda$ and non-conserved $T^{\mu\nu}$. The Bianchi identity imposes a relation between the energy-momentum (non)conservation and the variation of $G$ and $\Lambda$. If the energy-momentum $T^{\mu\nu}$ is conserved then both the Newton “constant” $G$ and the cosmological term $\Lambda$ either do not change or both must depend on time. If the energy-momentum $T^{\mu\nu}$ is not conserved then the Bianchi identity implies a relation between the energy-momentum and a variation in time of either $G$ or $\Lambda$ (or both). We apply thermodynamics in order to express the non-conservation of the energy-momentum of an ideal fluid by entropy and relate the time variations of $G$ and $\Lambda$ to a change of entropy. Using the relation between a varying Newton constant G and the black hole entropy we derive a modified formula for the Schwarzschild black hole evaporation (a slower evaporation). Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$. The average black hole’s density remains constant when the black hole’s radius shrinks to zero .
1 Introduction
We consider Einstein equations
$G^{\mu\nu}=\tilde{T}^{\mu\nu},$ (1)
where $G^{\mu\nu}=R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R$ $\mu,\nu=0,1,2,3$ and $R^{\mu\nu}$ is the Ricci tensor(we use the conventions of [1]). The form of the energy-momentum tensor $\tilde{T}^{\mu\nu}$ comes from a phenomenological description of fields and matter (we call them fluids). Then, the equations of motion for the fluids can be derived [2][3] from the covariant conservation law of the energy-momentum. In fact, in a wider perspective, the covariant conservation law for $\tilde{T}^{\mu\nu}$ can be applied for a description of fluid dynamics of systems (e.g.,quark-gluon plasma [4])in a flat space with a geometry of a submanifold embedded in the Minkowski space-time .
The standard Einstein equations can be expressed in the form (velocity of light $c=1$)
$G^{\mu\nu}=8\pi GT^{\mu\nu}-g^{\mu\nu}\Lambda.$ (2)
[…]
In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ of an ideal fluid. The violation of the energy-momentum conservation can be a consequence of an incomplete theory ( e.g. some particles (fields) are not taken into account). Some approximations (e.g., diffusion approximation for a particle motion) directly imply a dissipative dynamics. The energy-momentum conservation can be violated if quantum phenomena are included in classical theory. We suggest that the quantum thermodynamic energy conservation (the first law of thermodynamics) can be identified with the conservation of $\tilde{T}^{\mu\nu}$ if we admit a time dependence of $G$ or $\Lambda$ (or both) in the classical Einstein equations (2).
In sec.2 we obtain Bianchi identities resulting from a conservation of $\tilde{T}^{\mu\nu}$(1) and non-conservation of $T^{\mu\nu}$ (2). We show that the conservation of $T^{\mu\nu}$ allows a variation in time of both $G$ and $\Lambda$ but $\partial_{t}G\simeq\partial_{t}\Lambda$. Such models have been discussed in [17][18][19][20].These authors show that a variation of $G$ leads to a variation of $\Lambda$ in an agreement with observations of the universe expansion. In sec.3 applying the thermodynamics we express the variation of cosmological constants by the entropy $S$. According to the first law of thermodynamics $\nabla_{\mu}T^{\mu\nu}$ is related to the heat flow. Then, owing to the second law of thermodynamics the heat flow should be expressed by the change of entropy $TdS$ (where $T$ is the temperature). From the Bianchi identities (applied to eq.(2)) and the first law of thermodynamics it follows that a change of entropy is impossible if $G$ and $\Lambda$ are constant. We discuss a change of $G$ in models with an entropy generated by quantum phenomena at the event horizon. We apply the relation between the entropy and a varying Newton constant $G$ to the Schwarzschild black hole evaporation. As a consequence we obtain a modified evaporation formula. We obtain a slower black hole evaporation. Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$ (that the varying $G$ could change the evaporation law has been discussed in [11][12]). The black hole’s volume tends to zero, but the average energy density defined as a ratio of the mass by the volume does not change during the whole evaporation process ( in contradistinction with the case of a fixed Newton constant). The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [21][22] raised by an observation of a giant emission of neutrino radiation in KM3NeT [23] (for a search of a final burst of radiation from primordial black holes see [24] and references cited there).
2 The Bianchi identity
Einstein equations (2) can be derived from the action
$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big),$
where $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for an ideal fluid [25][26]. The Bianchi identity applied to eq.(1) leads to the equation
$\nabla_{\mu}\tilde{T}^{\mu}_{\nu}=0.$ (3)
It is understood that the solution of gravity in the presence of “fluids” means a solution of both eqs.(1) and (3).
The Bianchi identity (3) with $\tilde{T}^{\mu\nu}$ of eq.(2) gives an equation for the “fluids”
$8\pi\partial_{\mu}GT^{\mu}_{\nu}+8\pi G\nabla_{\mu}T^{\mu}_{\nu}-\partial_{\nu}\Lambda=0.$ (4)
We assume that the Riemannian manifold is globally hyperbolic . On a globally hyperbolic manifold one can introduce coordinates [27] such that the metric takes the form ($j,k=1,2,3$, we use the notation which will be applied in subsequent sections)
$dl^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-|g_{00}|dt^{2}+\gamma_{jk}dx^{j}dx^{k}.$ (5)
We assume that $G$ and $\Lambda$ depend only on time. Then, eq.(4) reads
$8\pi\partial_{t}GT^{0}_{\nu}+8\pi G\nabla_{\mu}T^{\mu}_{\nu}-\delta^{0}_{\nu}\partial_{t}\Lambda=0.$ (6)
At $\nu=0$ we have
$8\pi\partial_{t}GT^{0}_{0}+8\pi G\nabla_{\mu}T^{\mu}_{0}-\partial_{t}\Lambda=0.$ (7)
At the spatial index $j$
$\partial_{t}GT^{0}_{j}+G\nabla_{\mu}T^{\mu}_{j}=0.$ (8)
We assume the energy-momentum tensor $T^{\mu\nu}$ in the form of an ideal fluid
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p,$ (9)
where $\rho$ is the energy density and $p$ is the pressure. The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. For the ideal fluid we have
$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho.$ (10)
As an example of the metric (5) let us consider the FLRW metric describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
$dl^{2}=-dt^{2}+a^{2}\Big((1-kr^{2})^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\Big).$ (11)
In the metric (11) in the comoving frame $u=(1,{\bf 0})$ eq.(8) is satisfied if $p$ does not depend on spatial variables as $T^{0}_{j}=0$ and $\nabla_{j}T^{j}_{k}=\partial_{k}p=0$. The fluid velocity satisfies $\nabla_{\mu}u^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}u^{\mu})=3H$ with $H=a^{-1}\partial_{t}a$. Hence, if $\rho$ and $p$ depend only on time then from eqs.(7) and (9)
$8\pi\partial_{t}G\rho+8\pi G(\partial_{t}\rho+3H(\rho+p))+\partial_{t}\Lambda=0.$ (12)
If the energy-momentum tensor is conserved
$\partial_{t}\rho^{c}+3H(\rho^{c}+p^{c})=0$
then eq.(12) reads
$8\pi\partial_{t}G\rho^{c}=-\partial_{t}\Lambda.$ (13)
We obtain the standard formula for $\rho^{c}(a)$ from the energy-momentum conservation, but according to eq.(12) both $G$ and $\Lambda$ must depend on $a$. This is the case discussed by Gupta [17], Costa et al[18] and other authors [19] [20].
For the solution of eq.(13) it is useful to introduce the variable
$\tau=\ln a.$ (14)
We assume the linear equation of state $p^{c}=w^{c}\rho^{c}$ . From eq.(13) we can express $\Lambda$ by $G$ (we have $\rho^{c}=\rho^{c}_{0}a^{-3(1+w^{c})}$)
$\Lambda=\Lambda_{ex}-\rho^{c}_{0}\int_{\tau_{0}}^{\tau}8\pi\partial_{s}G\exp(-3(1+w^{c})s)ds,$ (15)
where $\Lambda_{ex}$ is an integration constant at $\tau_{0}$. We can interpret this formula as an appearance of a time-dependent cosmological term from a varying coupling constant [17][19].
In general, if the energy-momentum is not conserved, then we get from eq.(12)
$8\pi\partial_{t}(\rho G)+24\pi G\rho H(1+w)=-\partial_{t}\Lambda.$ (16)
By integration we can express $\Lambda$ in terms of $G\rho$ or $G\rho$ by $\Lambda$
$\begin{array}[]{l}\rho G=(\rho G)_{0}\exp(-3(1+w)(\tau-\tau_{0}))\cr-\frac{1}{8\pi}\int_{\tau_{0}}^{\tau}\exp(-3(1+w)(\tau-s))\partial_{s}\Lambda ds.\end{array}$ (17)
Let us note that choosing $\tau_{0}=$-$\infty$ (corresponding to the initial $a=0$) we get a solution with the initial $(\rho G)_{0}\exp(3(1+w)\tau_{0})=0$. If we assume $\Lambda=\Lambda_{ex}+\Lambda_{0}a^{\alpha}$, where $\Lambda_{0}$ is another constant, then in the Friedmann equation for the metric (11) we obtain from eq.(17) a modified matter distribution ( with $w_{\alpha}=-1-\frac{\alpha}{3}$; a phantom matter if $\alpha>0$ [6][28][29][30]).
3 Thermodynamics of the moving fluid
The first law of thermodynamics when applied to the energy-momentum $\tilde{T}^{\mu\nu}$ (1) means that the supply of heat $Q$ to the system in a volume $V$ yields a work $pdV$ done upon the system plus a change of its internal energy $E$
$dE+pdV=dQ=TdS.$ (18)
On the rhs of eq.(18) we have used the second law of thermodynamics (defining the entropy $S$) saying that there exists an integrating factor (the absolute temperature $T$) such that $dS=T^{-1}dQ$ depends only on the state of the system.
We are going to generalize eq.(18) so that it describes thermodynamics of an ideal fluid moving with a velocity $u^{\mu}$ in a metric more general than in eq.(11). Let us introduce the inverse temperature $\beta$ as a four-vector (for some recent discussions of an inverse four-temperature and the four-entropy, see [31][32][33] and references cited there)
$\beta^{\mu}=\beta u^{\mu}.$ (19)
The zeroth component
$\beta^{0}=\beta u^{0}=\beta|g_{00}|^{-\frac{1}{2}}(1+g_{jk}u^{j}u^{k})^{\frac{1}{2}}$ (20)
in the comoving frame $u=(u^{0},{\bf 0})$ is the Tolman-Ehrenfest equilibrium temperature [34], $T^{0}=T\sqrt{|g_{00}|}$, where $T=\beta^{-1}$ and $T^{0}=(\beta^{0})^{-1}$ (we choose units with the Boltzmann constant $k_{B}=1$). Then, we introduce the entropy density $\sigma^{\mu}$ as the four-vector
$\sigma^{\mu}=\sigma u^{\mu}.$ (21)
We express the first law of thermodynamics of a moving fluid as
$\beta^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-\nabla_{\mu}\sigma^{\mu}.$ (22)
We multiply the Bianchi identity (4) by $\beta^{\nu}$
$8\pi\partial_{\mu}G\beta^{\nu}T^{\mu}_{\nu}+8\pi G\beta^{\nu}\nabla_{\mu}T^{\mu}_{\nu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (23)
Then, inserting (22) in eq.(23) we obtain
$8\pi\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-8\pi G\nabla_{\mu}\sigma^{\mu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (24)
This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon. In the example of the metric (11) we have (in the comoving coordinates)
$\begin{array}[]{l}\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}\cr=u^{\mu}\nabla_{\mu}\sigma+3H\sigma=\partial_{t}\sigma+3H\sigma.\end{array}$ (25)
Hence, eq.(24) reads
$8\pi\beta\rho\partial_{t}G+\beta\partial_{t}\Lambda=-8\pi G(\partial_{t}\sigma+3H\sigma)$ (26)
or if we apply eq.(12)
$\partial_{t}\rho+3H{\cal G}=T\partial_{t}\sigma.$ (27)
where ${\cal G}=\rho+p-\sigma T$ is the density of the Gibbs potential. If we set $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ in eq.(27) then we recover eq.(18).
4 Compact systems with an event horizon
Eq.(24) may have a wider application to space-time phenomena. We have in mind the star contraction [35] or universe evolution ( e.g.,for time-dependent metrics relevant for black holes in an expanding universe [36][37][38]) . In this section we apply an integrated form of eq.(24) to a compact object. We integrate eq.(24) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric in eq.(5) ( for a general metric a better way would be to do the space-time integral in the time-interval $(t,t+dt)$ as in [39])
$\int d{\bf x}\sqrt{g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (28)
We assume that $G$ and $\Lambda$ depend only on time. In the comoving frame from eq.(20) $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$. Calculating the integral (28) of the energy-momentum of an ideal fluid and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta\rho+G\partial_{t}\int d{\bf x}\sqrt{g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (29)
We may skip the term $G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})$ as in the comoving frame $\sigma^{j}=\sigma u^{j}=0$. Denoting
$S=\int d{\bf x}\sqrt{\gamma}\sigma$ (30)
using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(29) as
$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}\rho+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (31)
We define the mass
$M=\int d{\bf x}\sqrt{\gamma}\rho.$ (32)
We skip the time derivative of $\Lambda$ in eq.(31)for an application to the black hole. Then, eq.(31) can be rewritten as
$M\partial_{t}G=-GT\partial_{t}S.$ (33)
We intend to apply the formulae (2) and (24) to gravitational systems with an event horizon whose entropy and temperature has been determined by other methods [40][41][42]. In classical closed systems the total entropy does not change ( neither does it in quantum systems). In systems with an event horizon we may restrict ourselves to subsystems located inside or outside the horizon. Then, their entropy may depend on time. The corresponding calculations have been performed in quantum systems ( calculation of the entanglement entropy) [43][44][45]( see the reviews in [46][47]). We think that this entropy can be incorporated in eq.(33). Eq.(33) is a consequence of eq.(24). In eq.(24) we have replaced the covariant derivative of the energy-momentum tensor (the lhs of eq.(18)) appearing in the Bianchi identity (23) by the covariant derivative of the entropy density according to the first law of thermodynamics (the rhs of eq. (18)). The quantum results for the entropy and for the temperature (obtained by calculations of the partition function [43] or from the quantum gravity in the Euclidean formulation [42]) show that $S\simeq\frac{1}{\hbar}$ and $T$ is proportional to $\hbar$. Hence, their product on the rhs of eqs.(27) and (33) has a classical meaning.
The formula for the temperature of the Schwarzschild black hole of mass $M$ is [41]
$T_{BH}=\frac{\hbar}{8\pi GM}.$ (34)
It can be derived from the correlation functions of the quantum scalar fields on the Schwarzschild background or from the Rindler space approximation on the basis of the KMS periodicity argument. Another derivation is based on Euclidean quantum gravity in one-loop approximation [42]. The entropy of the Schwarzschild black hole can be calculated from the second law of thermodynamics $dM=T_{BH}dS_{BH}$ on the basis of eq.(34)
$S_{BH}=\frac{4\pi}{\hbar}M^{2}G$ (35)
(it is also obtained from a path integral as an entanglement entropy [42][43][44][45], in such a case it would appear in the semi-classical expansion of the effective action). We insert the black hole entropy (35) in eq.(33). Then, we obtain
$\frac{3}{2}\frac{d\ln G}{dt}=-\frac{d\ln M}{dt}.$ (36)
Hence,
$G=G_{0}M_{0}^{\frac{2}{3}}M^{-\frac{2}{3}}.$ (37)
The change of mass $M$ resulting from the black hole evaporation through an emission of massless particles is described by the formula [48] [49][50][51] (see also the reviews in [46][47])
$\frac{dM}{dt}=-\hbar\alpha M^{-2}G^{-2},$ (38)
where $\alpha$ is a number which can be obtained from Stefan-Boltzmann formula multiplied by the black hole surface and an estimate of emissivity [48] ( the formula does not apply if $T_{BH}$ is lower than the temperature of the environment and for a small time). Eqs.(38) and (33) express the thermodynamic conservation law. The black hole is releasing the energy and the entropy gained during a mass accretion in the process of black hole formation.
Eq.(38) has the solution
$M^{3}=M_{0}^{3}-3\hbar\alpha\int_{t_{0}}^{t}G^{-2}(s)ds,$ (39)
where $t_{0}$ is the time of the black hole formation. We can insert this $M$ in eq.(37) for $G$. Denote
$X=\int_{t_{0}}^{t}G^{-2}(s)ds.$ (40)
Then, eq.(37) can be rewritten as
$\frac{dX}{dt}=G_{0}^{-2}M_{0}^{-\frac{4}{3}}(M_{0}^{3}-3\hbar\alpha X)^{\frac{4}{9}}.$ (41)
The resulting equation for $G$ has the solution
$G=G_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{2}{5}}.$ (42)
Inserting $G$ of eq.(42) into the formula (39) for the mass we obtain
$M=M_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{3}{5}}\equiv M_{0}Z^{\frac{3}{5}}.$ (43)
It follows that
$\begin{array}[]{l}T_{BH}=\frac{\hbar}{8\pi G_{0}M_{0}}Z^{-\frac{1}{5}}\end{array}$ (44)
The entropy tends to zero as $GM^{2}\simeq Z^{\frac{4}{5}}$, the luminosity $I$ of the black hole is proportional to $T_{BH}^{4}$ and to the area of the black hole ($A\simeq(GM)^{2}$ ). Hence, the luminosity behaves as
$I\simeq Z^{-\frac{2}{5}}$ (45)
tending to infinity at the end of the black hole’s life time with a speed lower than the one for a fixed Newton constant.
The black hole’s radius $2GM$ tends to zero as $Z^{\frac{1}{5}}$ but the average energy density defined as the mass divided by the volume remains constant as
$M\Big(\frac{4\pi}{3}(2GM)^{3}\Big)^{-1}=M_{0}\Big(\frac{4\pi}{3}(2G_{0}M_{0})^{3}\Big)^{-1}$ (46)
This is different than in the Hawking-Page model [52][51] when the averaged energy density tends to infinity at the end of black hole’s life time. The results (42)-(46) could be compared with observations of the gamma-ray bursts [24]. In particular, with the recent observations of the giant neutrino emissions [23] which by some authors [21][22] are interpreted as resulting from primordial black hole explosions.
The life time of the black hole according to eq.(43) is
$t_{f}=\frac{3}{5}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (47)
It is longer by a factor $\frac{9}{5}$ than the one of the Hawking-Page model (which would result from the assumption that $G=const$ in eq.(39)).
5 Summary and outlook
We have shown that the Bianchi identity of classical Einstein equations admits a violation of the ideal fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time. As a generalization of the first and the second laws of thermodynamics the covariant energy-momentum divergence can be expressed by the covariant divergence of the entropy density. Then, the entropy density enters the Bianchi identity. In this paper we have applied the formula for the change of entropy during the black hole evaporation to determine the time-dependence of $G$ in the Bianchi identity. As a consequence of a growing $G$ the black hole evaporation is slower and the average black hole density is constant in contradistinction to the one for a fixed $G$ (this could be subject to astronomical observations) . The time-dependence of $G$ is unavoidable if we wish to describe the non-conserved energy-momentum of an evaporating black hole by Einstein equations (2). The increasing $G$ could have a physical stabilizing origin preventing the appearance of an infinite density when the black hole’s radius tends to zero. Another problem deserving a study concerns the entropy of an expanding universe with a cosmological horizon. For such a system the varying cosmological constant in the Bianchi identity should be relevant.
Data availability statement No data were created or analysed in this study
Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis. JH thanks Dr. Aleksander Kozak for supervising the thesis.

### v2

@@
 Abstract
-In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of an ideal fluid plus the cosmological term.
+In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of a fluid plus the cosmological term.
 We consider time-dependent Newton “constant” $G$, the cosmological term $\Lambda$ and non-conserved $T^{\mu\nu}$.
 The Bianchi identity imposes a relation between the energy-momentum (non)conservation and the variation of $G$ and $\Lambda$.
-If the energy-momentum $T^{\mu\nu}$ is conserved then both the Newton “constant” $G$ and the cosmological term $\Lambda$ either do not change or both must depend on time.
-If the energy-momentum $T^{\mu\nu}$ is not conserved then the Bianchi identity implies a relation between the energy-momentum and a variation in time of either $G$ or $\Lambda$ (or both).
-We apply thermodynamics in order to express the non-conservation of the energy-momentum of an ideal fluid by entropy and relate the time variations of $G$ and $\Lambda$ to a change of entropy.
-Using the relation between a varying Newton constant G and the black hole entropy we derive a modified formula for the Schwarzschild black hole evaporation (a slower evaporation).
-Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$.
-The average black hole’s density remains constant when the black hole’s radius shrinks to zero .
+The covariant divergence $\nabla_{\mu}T^{\mu\nu}$ can be related to the first law of thermodynamics.
+For compact systems of mass $M$ from the Bianchi identity we obtain a power-law relation $G\simeq M^{-\gamma}$ with $\gamma$ depending on pressure or entropy.
+We discuss radiation and a mass loss described by the Stefan-Boltzmann law.
+In this formula we insert an expression for the black hole area and its temperature $T$.
+The Bianchi identity together with a formula for temperature and entropy $S$ determines the index $\gamma$ in the relation between the Newton constant $G$ and the mass $M$.
+If the entropy $S$ is defined by the equation $dS=T^{-1}dM$ then $\gamma=1$ (the same as for zero pressure).
+If the formula of Bekenstein-Hawking entropy holds true for time-dependent $G$ then $\gamma=\frac{2}{3}$.
+We discuss consequences for the evaporation law of some modified expressions for the entropy appearing in effective models of gravity resulting from an interaction with matter fields.
+In particular, $\gamma=1$ leads to a constant evaporation temperature whereas $\gamma>1$ to a decreasing temperature and luminosity.
 1 Introduction
@@
 Then, the equations of motion for the fluids can be derived [2][3] from the covariant conservation law of the energy-momentum.
-In fact, in a wider perspective, the covariant conservation law for $\tilde{T}^{\mu\nu}$ can be applied for a description of fluid dynamics of systems (e.g.,quark-gluon plasma [4])in a flat space with a geometry of a submanifold embedded in the Minkowski space-time .
 The standard Einstein equations can be expressed in the form (velocity of light $c=1$)
@@
 […]
-In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ of an ideal fluid.
+In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ .
 The violation of the energy-momentum conservation can be a consequence of an incomplete theory ( e.g.
@@
 Such models have been discussed in [19][20][21][22].These authors show that a variation of $G$ leads to a variation of $\Lambda$ in an agreement with observations of the universe expansion.
-In sec.3 applying the thermodynamics we express the variation of cosmological constants by the entropy $S$.
-According to the first law of thermodynamics $\nabla_{\mu}T^{\mu\nu}$ is related to the heat flow.
-Then, owing to the second law of thermodynamics the heat flow should be expressed by the change of entropy $TdS$ (where $T$ is the temperature).
+In secs.3 and 4 we show that $\nabla_{\mu}T^{\mu\nu}$ can be related to the heat $Q$ according to the formula $dQ=TdS$ where $T$ is the temperature and $S$ is the entropy.
+We apply this formula to non-equilibrium systems with time-dependent variables.
 From the Bianchi identities (applied to eq.(2)) and the first law of thermodynamics it follows that a change of entropy is impossible if $G$ and $\Lambda$ are constant.
-We discuss a change of $G$ in models with an entropy generated by quantum phenomena at the event horizon.
-We apply the relation between the entropy and a varying Newton constant $G$ to the Schwarzschild black hole evaporation.
-As a consequence we obtain a modified evaporation formula.
-We obtain a slower black hole evaporation.
-Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$ (that the varying $G$ could change the evaporation law has been discussed in [11][12]).
-The black hole’s volume tends to zero, but the average energy density defined as a ratio of the mass by the volume does not change during the whole evaporation process ( in contradistinction with the case of a fixed Newton constant).
-The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [21][22] raised by an observation of a giant emission of neutrino radiation in KM3NeT [23] (for a search of a final burst of radiation from primordial black holes see [24] and references cited there).
+We obtain an equation which relates the time derivatives of $G$ and the mass $M$ of a compact system.
+If the pressure in the system is zero then this equation implies that $GM$ is constant.
+In some systems the Bianchi identity leads to a power-law relation $G\simeq M^{-\gamma}$.
+In sec.5 we discuss a compact system with an event horizon and its mass decrease resulting from a radiation described by the Stefan-Boltzmann law.
+In such systems the temperature as well as the area of the horizon depend on $G$ and $M$.
+When the Bianchi identity determines the relation between $G$ and $M$ then the Stefan-Boltzmann law predicts an explicit dependence on time of both $G$ and $M$.
+We insert the Bekenstein-Hawking formula for the temperature in this radiation law.
+In a system with zero pressure as a consequence of a constant $GM$ we obtain constant temperature and constant luminosity during the whole evaporation process (no black hole explosion).
+We derive the same conclusion from the first law of thermodynamics if the entropy is determined by the equation $dS=T^{-1}dM$.
+We obtain a different result if we assume that Bekenstein-Hawking formula for the entropy applies without any change to models with a time-dependent $G$.
+In such a case the Bekenstein-Hawking temperature as well as the black hole luminosity tend to infinity at the end of the evaporation process but the mean black hole density remains constant.
+We discuss the evaporation formula for various possible $\gamma$ in some other models of gravity.
+In particular, $\gamma>1$ leads to a decreasing black hole temperature and a decreasing luminosity (that the varying $G$ could change the evaporation law has been discussed in [10][11]).
+The compact systems with an event horizon could lead to useful tests (besides the ones discussed in [17][18]) of a varying Newton constant $G$ because the area and the temperature of the horizon depend on $G$.
+The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [23][24] raised by an observation of a giant emission of neutrino radiation in KM3NeT [25] (for a search of radiation from primordial black holes see [26] and references cited there).
+A decreasing black hole temperature would be welcome in models of primordial black holes as a cold dark matter.
 2 The Bianchi identity
 Einstein equations (2) can be derived from the action
-$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big),$
-where $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for an ideal fluid [25][26]. The Bianchi identity applied to eq.(1) leads to the equation
+$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big)+K_{GHY},$
+where $K_{GHY}$ is the Gibbons-Hawking-York surface term [27][28], $x=(t,{\bf x})$, $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for a fluid [29][30].
+The Bianchi identity applied to eq.(1) leads to the equation
 $\nabla_{\mu}\tilde{T}^{\mu}_{\nu}=0.$ (3)
@@
 $\partial_{t}GT^{0}_{j}+G\nabla_{\mu}T^{\mu}_{j}=0.$ (8)
-We assume the energy-momentum tensor $T^{\mu\nu}$ in the form of an ideal fluid
-$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p,$ (9)
-where $\rho$ is the energy density and $p$ is the pressure. The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. For the ideal fluid we have
-$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho.$ (10)
-As an example of the metric (5) let us consider the FLRW metric describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
+We consider the energy-momentum tensor $T^{\mu\nu}$ of a fluid
+$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p+\Sigma^{\mu\nu},$ (9)
+where $\rho$ is the energy density, $p$ is the pressure and $\Sigma^{\mu\nu}$ describes a dissipation (“viscosity” ) of the fluid [32][33] .
+The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$.
+We have
+$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho+u^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu}.$ (10)
+As an example of the metric (5) let us consider the FLRW solution describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
 $dl^{2}=-dt^{2}+a^{2}\Big((1-kr^{2})^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\Big).$ (11)
+The metric (11) is a solution of Einstein equations with an ideal homogeneous fluid ($\Sigma=0$).
 In the metric (11) in the comoving frame $u=(1,{\bf 0})$ eq.(8) is satisfied if $p$ does not depend on spatial variables as $T^{0}_{j}=0$ and $\nabla_{j}T^{j}_{k}=\partial_{k}p=0$.
@@
 $8\pi\partial_{t}G\rho+8\pi G(\partial_{t}\rho+3H(\rho+p))+\partial_{t}\Lambda=0.$ (12)
+A generalization of eqs.(11)-(12) to $\Sigma\neq 0$ is interesting because it may describe a pasting of FLWR solution with the Schwarzschild solution describing a black hole in an expanding metric [34][35][36][37].
 If the energy-momentum tensor is conserved
@@
 3 Thermodynamics of the moving fluid
+We multiply the Bianchi identity (4) by the velocity $u$, integrate with $\sqrt{-g}d{\bf x}$ and write it in a form showing a relation to the first law of thermodynamics
+$\int d{\bf x}\sqrt{-g}\Big(u^{\mu}\partial_{\mu}(G\rho)+G(\rho+p)\nabla_{\mu}u^{\mu}+u^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu})\Big)=-\frac{1}{8\pi}\int d{\bf x}\sqrt{-g}u^{\mu}\partial_{\mu}\Lambda,$ (18)
+where eq.(10) has been applied.
+We express eq.(18) in the comoving frame in the metric (5) ( then $g_{00}u^{0}u^{0}=-1$ and $\sqrt{-g}=\sqrt{-g_{00}}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric).
+We use
+$\nabla_{\mu}u^{\mu}=(\sqrt{-g})^{-1}\partial_{\mu}(u^{\mu}\sqrt{-g})=(\sqrt{-g})^{-1}\partial_{t}\sqrt{\gamma}.$ (19)
+Then, eq.(18) reads
+$\begin{array}[]{l}\int d{\bf x}\sqrt{\gamma}(\partial_{t}G\rho+G\partial_{t}\rho)+G\int d{\bf x}(\rho+p)\partial_{t}\sqrt{\gamma}\cr=-\int d{\bf x}\sqrt{\gamma}\nabla_{t}\Sigma^{0}_{0}-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (20)
 The first law of thermodynamics when applied to the energy-momentum $\tilde{T}^{\mu\nu}$ (1) means that the supply of heat $Q$ to the system in a volume $V$ yields a work $pdV$ done upon the system plus a change of its internal energy $E$
@@
 On the rhs of eq.(21) we have used the second law of thermodynamics (defining the entropy $S$) saying that there exists an integrating factor (the absolute temperature $T$) such that $dS=T^{-1}dQ$ depends only on the state of the system.
-We are going to generalize eq.(18) so that it describes thermodynamics of an ideal fluid moving with a velocity $u^{\mu}$ in a metric more general than in eq.(11).
+Eq.(21) can be rewritten in the form
+$d{\cal H}-Vdp=TdS,$ (22)
+where ${\cal H}=E+pV$ is the enthalpy.
+It can be seen that the lhs of eq.(20) coincides with the lhs of eq.(21) if we extend the meaning of the differentials of neighboring configurations to configurations at different times.
+The term $\Sigma$ related to the heat flow in phenomenological thermodynamics is usually also expressed by entropy.
+We shall assume that it is already contained in $T\partial_{t}S$ and ignore it.
+Note that if $p=0$ then eq.(20) could be interpreted as the formula (22) for enthalpy with $p$ replaced by $\frac{1}{8\pi}\Lambda$ in the formulation of the black hole pressure according to to ref.[41].
+When we define the mass as
+$M=\int d{\bf x}\sqrt{\gamma}\rho$ (23)
+then eqs.(20)-(21) give
+$\begin{array}[]{l}M\partial_{t}G+G(\partial_{t}M+\int d{\bf x}p\partial_{t}\sqrt{\gamma})=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda\end{array}$ (24)
+or in terms of the entropy
+$\begin{array}[]{l}M\partial_{t}G+GT\partial_{t}S=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (25)
+In eqs.(23) and (25) we introduced the volume (for the metric (5))
+$dV=d{\bf x}\sqrt{\gamma}$ (26)
+with $d\partial_{t}V=d{\bf x}\partial_{t}\sqrt{\gamma}$.
+For $p=0$ eq.(24) takes a simple form which will be applied in sec.5
+$\begin{array}[]{l}\partial_{t}(MG)=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (27)
+There is a long-standing debate over the notion of the energy, volume and the mass in the context of thermodynamics in general relativity [42] (for a recent discussion see [43]).
+In the classic paper [44] the mass has the ADM meaning [1] of the mass measured at infinity and evaluated on the basis of its gravitational effects.
+The mass (23) is the mass of the fluid.
+For relativistic stars eq.(23) agrees with Weinberg’s definition of the mass ( ref.[45], sec.11, eq.(11.1.23)) if $\rho$ is interpreted as the density of nucleons in the star.
+The mass (23) appears in the formulas for the neutron star.
+It leads to the relation $G\simeq M^{-\frac{2}{3}}$ ([45], sec.11)in the formula for the Chandrasekhar limt of the neutron star.
+According to ref.[18] it gives a bound on the variation of $G$.
+In geometries pasting the FLWR geometry (5) with the Schwarzschild geometry (as discussed in sec.5) these various definitions of the mass may be related.
+The reason that we integrated the covariant equation (4) with $d{\bf x}\sqrt{-g}$ is that the space-time volume $dtd{\bf x}\sqrt{-g}$ in eq.(18) does not depend on the choice of coordinates.
+We could integrate in eq.(18) with $dtd{\bf x}\sqrt{-g}v$ with an arbitrary function $v$.
+Then, in the integrated form of eq.(24) ( independent of the choice of coordinates if $v$ is a scalar)
+$\begin{array}[]{l}\int dt\Big(\tilde{M}\partial_{t}G+G\partial_{t}(\tilde{M}+G\int d{\bf x}vp\partial_{t}\sqrt{\gamma}\Big)\cr=-\frac{1}{8\pi}\int dt\int d{\bf x}\sqrt{\gamma}v\partial_{t}\Lambda\end{array}$
+with $\tilde{M}=\int d{\bf x}\sqrt{\gamma}v\rho$.
+The only change would involve a change of measure $\sqrt{\gamma}\rightarrow v\sqrt{\gamma}$ in the definition of the mass in eq.(23) and the entropy in terms of an entropy density in eq.(25).
+We suggest a generalization of eq.(20) to an explicitly covariant formula for relativistic fluids.
 Let us introduce the inverse temperature $\beta$ as a four-vector (for some recent discussions of an inverse four-temperature and the four-entropy, see [33][46][47] and references cited there)
@@
 $8\pi\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-8\pi G\nabla_{\mu}\sigma^{\mu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (33)
-This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon. In the example of the metric (11) we have (in the comoving coordinates)
-$\begin{array}[]{l}\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}\cr=u^{\mu}\nabla_{\mu}\sigma+3H\sigma=\partial_{t}\sigma+3H\sigma.\end{array}$ (25)
-Hence, eq.(24) reads
-$8\pi\beta\rho\partial_{t}G+\beta\partial_{t}\Lambda=-8\pi G(\partial_{t}\sigma+3H\sigma)$ (26)
-or if we apply eq.(12)
-$\partial_{t}\rho+3H{\cal G}=T\partial_{t}\sigma.$ (27)
-where ${\cal G}=\rho+p-\sigma T$ is the density of the Gibbs potential. If we set $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ in eq.(27) then we recover eq.(18).
-4 Compact systems with an event horizon
-Eq.(24) may have a wider application to space-time phenomena.
-We have in mind the star contraction [35] or universe evolution ( e.g.,for time-dependent metrics relevant for black holes in an expanding universe [36][37][38]) .
-In this section we apply an integrated form of eq.(24) to a compact object.
-We integrate eq.(24) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric in eq.(5) ( for a general metric a better way would be to do the space-time integral in the time-interval $(t,t+dt)$ as in [39])
-$\int d{\bf x}\sqrt{g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (28)
+This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon.
+We have $\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}$.
+Hence, if we use (4) and (10) then eq.(31) can be rewritten in the form
+$u^{\mu}\partial_{\mu}\rho+(\nabla_{\mu}u^{\mu}){\cal G}=\beta^{-1}u^{\mu}\partial_{\mu}\sigma-\beta^{-1}\beta^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu}.$ (34)
+where ${\cal G}=\rho+p-\beta^{-1}\sigma$ is the density of the Gibbs potential.
+As (in the comoving frame) $\nabla_{\mu}u^{\mu}=(\sqrt{\gamma})^{-1}\partial_{t}\sqrt{\gamma}\equiv V^{-1}\partial_{t}V$ then (ignoring $\Sigma$)
+$\partial_{t}\rho+V^{-1}\partial_{t}V{\cal G}=T\partial_{t}\sigma.$ (35)
+When we insert in eq.(35) $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ then we recover eq.(21). We can also rewrite eq.(35) in the form
+$\beta^{-1}\partial_{t}(V\sigma)=\partial_{t}(V\rho)+p\partial_{t}V$ (36)
+leading to the correct formula for static systems $\sigma=\beta(\rho+p)$.
+4 Compact systems
+In this section we show that eq.(31) leads to the same conclusions as eqs.(24)-(25) and (27) of sec.3. We integrate eq.(33) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$
+$\int d{\bf x}\sqrt{-g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (37)
 We assume that $G$ and $\Lambda$ depend only on time.
-In the comoving frame from eq.(20) $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$.
-Calculating the integral (28) of the energy-momentum of an ideal fluid and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
-$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta\rho+G\partial_{t}\int d{\bf x}\sqrt{g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (29)
+In the comoving frame $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$ from eq.(28).
+Calculating the integral (37) of the energy-momentum and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
+$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta T^{0}_{0}+G\partial_{t}\int d{\bf x}\sqrt{-g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{-g}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (38)
 We may skip the term $G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})$ as in the comoving frame $\sigma^{j}=\sigma u^{j}=0$. Denoting
 $S=\int d{\bf x}\sqrt{\gamma}\sigma$ (39)
-using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(29) as
-$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}\rho+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (31)
-We define the mass
-$M=\int d{\bf x}\sqrt{\gamma}\rho.$ (32)
-We skip the time derivative of $\Lambda$ in eq.(31)for an application to the black hole. Then, eq.(31) can be rewritten as
+using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(37) in the form (25)
+$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}T^{0}_{0}+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (40)
+When we skip the time derivative of $\Lambda$ in eq.(40)and denote $M=\int d{\bf x}\sqrt{\gamma}T^{0}_{0}$ (we admit that with $\Sigma^{0}_{0}\neq 0$ the spatial integrals over $\rho$ and $T^{0}_{0}$ may be different) then we obtain
 $M\partial_{t}G=-GT\partial_{t}S.$ (41)
-We intend to apply the formulae (2) and (24) to gravitational systems with an event horizon whose entropy and temperature has been determined by other methods [40][41][42].
-In classical closed systems the total entropy does not change ( neither does it in quantum systems).
-In systems with an event horizon we may restrict ourselves to subsystems located inside or outside the horizon.
-Then, their entropy may depend on time.
-The corresponding calculations have been performed in quantum systems ( calculation of the entanglement entropy) [43][44][45]( see the reviews in [46][47]).
-We think that this entropy can be incorporated in eq.(33).
-Eq.(33) is a consequence of eq.(24).
-In eq.(24) we have replaced the covariant derivative of the energy-momentum tensor (the lhs of eq.(18)) appearing in the Bianchi identity (23) by the covariant derivative of the entropy density according to the first law of thermodynamics (the rhs of eq.
-(18)).
-The quantum results for the entropy and for the temperature (obtained by calculations of the partition function [43] or from the quantum gravity in the Euclidean formulation [42]) show that $S\simeq\frac{1}{\hbar}$ and $T$ is proportional to $\hbar$.
-Hence, their product on the rhs of eqs.(27) and (33) has a classical meaning.
-The formula for the temperature of the Schwarzschild black hole of mass $M$ is [41]
-$T_{BH}=\frac{\hbar}{8\pi GM}.$ (34)
-It can be derived from the correlation functions of the quantum scalar fields on the Schwarzschild background or from the Rindler space approximation on the basis of the KMS periodicity argument.
-Another derivation is based on Euclidean quantum gravity in one-loop approximation [42].
-The entropy of the Schwarzschild black hole can be calculated from the second law of thermodynamics $dM=T_{BH}dS_{BH}$ on the basis of eq.(34)
+5 Mass loss in evaporation
+In this section we discuss dynamics of a change of mass of a compact object.
+We assume that such dynamics can be described in the framework of classical Einstein gravity if $G$ and $\Lambda$ are allowed to change in time.
+We have two equations (24) and (25) for a relation between time derivatives of $G$ and $M$.
+We have in mind a dynamical system resulting from a star contraction which subsequently is decaying through an emission of radiation.
+An investigation of the dynamics of a star collapse to the black hole has been initiated by Oppenheimer and Snyder [49].
+Then, Einstein and Straus [50] [51] studied a star (Schwarzschild metric) in an expanding universe by pasting the FLWR metric with the Schwarzschild solution.
+In this framework there is a lot of work recently on the formation of the black hole [52] [53], its evolution in the FLWR space-time [34][35][36][37] and final evaporation [54][55] (nonsingular black holes).
+In this section we apply eqs.(25) and (27) to a description of a decrease of mass as a result of an emission of radiation of a compact object (a black hole) with a hot event horizon.
+A change of mass $M$ resulting from an emission of radiation from a surface can be obtained from the Stefan-Boltzmann law (we assume that the temperature $T$ is larger then the temperature of the environment and that the absorption from the environment can be neglected)
+$\frac{dM}{dt}=-\lambda T^{4}A,$ (42)
+where $\lambda$ is the Stefan-Boltzmann constant ($\lambda\simeq\hbar^{-3}$) and $A$ is the area of the surface.
+For a compact system with an event horizon we can obtain a system of equations (24)-(25) and (42) depending on $G$ and $M$ because the area and temperature are functions of $G$ and $M$.
+In eq.(24) we would need to know the pressure whereas in eq.(25) the entropy as functions of $G$ and $M$ in order to determine a relation between these variables.
+We consider a power-law Ansatz as a possible solution of eqs.(24)-(25)
+$G=G_{0}M_{0}^{\gamma}M^{-\gamma}.$ (43)
+It can be seen that if $\partial_{t}\Lambda=0$ and $p=0$ then from eq.(24) $\gamma=1$.
+In order to determine a relation between $G$ and $M$ from eq.(25) we need to know the temperature $T$ and entropy $S$.
+We consider a simple power-law Ansatz
+$T=\theta G^{\alpha_{1}}M^{\alpha_{2}},$ (44)
+$S=\kappa G^{\nu_{1}}M^{\nu_{2}}.$ (45)
+We substitute it in eq.(25). Then, we obtain
+$\gamma=\theta\kappa\nu_{2}(1+\theta\kappa\nu_{1})^{-1}$ (46)
+with $\alpha_{1}=-\nu_{1}$$\alpha_{2}=1-\nu_{2}$. If in addition we demand the definition of temperature as
+$\frac{1}{T}=\frac{\partial S}{\partial M}$ (47)
+then we get the relation $\kappa\theta\nu_{2}=1$ and $\gamma=\nu_{2}(\nu_{1}+\nu_{2})^{-1}$.
+As conjectured by Bekenstein [56] the entropy of the black hole is proportional to its area.
+The calculation of the temperature of the black hole in a system of quantum scalar particles on a Schwarzschild background [57] gave the result (the formula has been established for various quantum systems defined on the Schwarzschild background [58][59])
+$T_{BH}=\frac{\hbar}{8\pi GM}$ (48)
+Then, from the relations (44)-(47) we obtain the entropy
 $S_{BH}=\frac{4\pi}{\hbar}M^{2}G$ (49)
-(it is also obtained from a path integral as an entanglement entropy [42][43][44][45], in such a case it would appear in the semi-classical expansion of the effective action).
-We insert the black hole entropy (35) in eq.(33).
-Then, we obtain
-$\frac{3}{2}\frac{d\ln G}{dt}=-\frac{d\ln M}{dt}.$ (36)
-Hence,
-$G=G_{0}M_{0}^{\frac{2}{3}}M^{-\frac{2}{3}}.$ (37)
-The change of mass $M$ resulting from the black hole evaporation through an emission of massless particles is described by the formula [48] [49][50][51] (see also the reviews in [46][47])
+If instead of (47) we use (48) and
+$dS=T_{BH}^{-1}dM$ (50)
+appearing in the black hole mechanics in [44] then the formula for the entropy is
+$S=\frac{8\pi}{\hbar}\int GMdM.$ (51)
+For a time-dependent $G$ eqs.(49) and (51) are different.
+We discuss all options (49), (51)and (44)-(45) leading to the relation (43) with different $\gamma$.
+The reason is that a time-dependent $G$ may appear in an effective field theory of various origin [60]: Higgs gauge theories with symmetry breaking [61][62],Branse-Dicke [10][11][63], $f(R)$ and higher derivatives gravities [64][63].
+For a constant $G$ the entropy (49) has been derived as an entanglement entropy in various models of quantum field theory [58][60][65][66].
+A time dependent temperature (possibly in eq.(51)) can appear if the black hole surface is time dependent ( a trapping horizon) in an expanding universe [67].
+A renormalization of the Newton constant as in refs.
+[13][14][15][16] can lead to a modified formula for the entropy.
+A microscopic counting of states in loop quantum gravity [68] does not fix the coefficient for entropy in front of the area as well.
+We consider also the option
+$S_{\kappa}=\frac{\kappa}{4G\hbar}A,$ (52)
+where $\kappa=G_{eff}^{-1}G$ with an effective coupling $G_{eff}$ which can come from a calculation of the entropy from the partition function.
+With the entropy (52) we get the relation (43) with $\gamma=\frac{\kappa}{1+\frac{1}{2}\kappa}$ .
+Summarizing, $\gamma=1$ for zero pressure in eq.(25) or the entropy (51), $\gamma=\frac{2}{3}$ for the Bekenstein-Hawking entropy (49).
+The remaining values of $\gamma\geq 0$ are rather hypothetical.
+However, they lead to interesting consequences.
+A constant $G$ (corresponding to $\gamma=0$ ) cannot describe evaporation by means of Einstein equations because owing to Bianchi identities decreasing $M$ (with constant $\Lambda$) requires a change of $G$.
+We assume in all models $T\simeq(GM)^{-1}$ and $A\simeq(GM)^{2}$. When we insert the temperature and the area in eq.(42) then we obtain [69] [70][71][72] (see also the reviews in [73][74][75])
 $\frac{dM}{dt}=-\hbar\alpha M^{-2}G^{-2},$ (53)
-where $\alpha$ is a number which can be obtained from Stefan-Boltzmann formula multiplied by the black hole surface and an estimate of emissivity [48] ( the formula does not apply if $T_{BH}$ is lower than the temperature of the environment and for a small time).
-Eqs.(38) and (33) express the thermodynamic conservation law.
-The black hole is releasing the energy and the entropy gained during a mass accretion in the process of black hole formation.
-Eq.(38) has the solution
-$M^{3}=M_{0}^{3}-3\hbar\alpha\int_{t_{0}}^{t}G^{-2}(s)ds,$ (39)
-where $t_{0}$ is the time of the black hole formation. We can insert this $M$ in eq.(37) for $G$. Denote
-$X=\int_{t_{0}}^{t}G^{-2}(s)ds.$ (40)
-Then, eq.(37) can be rewritten as
-$\frac{dX}{dt}=G_{0}^{-2}M_{0}^{-\frac{4}{3}}(M_{0}^{3}-3\hbar\alpha X)^{\frac{4}{9}}.$ (41)
-The resulting equation for $G$ has the solution
-$G=G_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{2}{5}}.$ (42)
-Inserting $G$ of eq.(42) into the formula (39) for the mass we obtain
-$M=M_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{3}{5}}\equiv M_{0}Z^{\frac{3}{5}}.$ (43)
-It follows that
-$\begin{array}[]{l}T_{BH}=\frac{\hbar}{8\pi G_{0}M_{0}}Z^{-\frac{1}{5}}\end{array}$ (44)
-The entropy tends to zero as $GM^{2}\simeq Z^{\frac{4}{5}}$, the luminosity $I$ of the black hole is proportional to $T_{BH}^{4}$ and to the area of the black hole ($A\simeq(GM)^{2}$ ).
-Hence, the luminosity behaves as
-$I\simeq Z^{-\frac{2}{5}}$ (45)
-tending to infinity at the end of the black hole’s life time with a speed lower than the one for a fixed Newton constant.
-The black hole’s radius $2GM$ tends to zero as $Z^{\frac{1}{5}}$ but the average energy density defined as the mass divided by the volume remains constant as
-$M\Big(\frac{4\pi}{3}(2GM)^{3}\Big)^{-1}=M_{0}\Big(\frac{4\pi}{3}(2G_{0}M_{0})^{3}\Big)^{-1}$ (46)
-This is different than in the Hawking-Page model [52][51] when the averaged energy density tends to infinity at the end of black hole’s life time.
-The results (42)-(46) could be compared with observations of the gamma-ray bursts [24].
+where $\alpha$ is a ($\hbar$ independent) number which can be obtained from the Stefan-Boltzmann constant and an estimate of emissivity [69].
+If the pressure is zero in eq.(24) (as in the Oppenheimer-Snyder approximation) leading to eq.(27) and $\partial_{t}\Lambda$ is negligible in eq.(27) then (as $GM\simeq const$) we obtain from eq.(53)
+$M_{t}=M_{0}-\hbar\alpha M_{0}^{-2}G_{0}^{-2}(t-t_{0}),$ (54)
+where $t_{0}$ is the time of the black hole formation. We obtain the same result ($\gamma=1$) from eq.(25) with an application of the formulas (50)-(51) for the entropy.
+With a varying $\Lambda$ and $\gamma=1$ we substitute from eq.(27) in eq.(53)
+$\begin{array}[]{l}MG=G_{0}M_{0}-\int_{t_{0}}^{t}\frac{ds}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{s}\Lambda.\end{array}$ (55)
+Then,
+$M=M_{0}-\hbar\alpha\int_{t_{0}}^{t}d\tau\Big(G_{0}M_{0}-\int_{t_{0}}^{\tau}\frac{ds}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{s}\Lambda\Big)^{-2}$ (56)
+Note that from eq.(55)it follows that if $\partial_{t}\Lambda=0$ then the temperature of the black hole as well as its luminosity are constant during the whole black hole’s life time (there is no black hole explosion, contrary to [72]).
+With the varying cosmological term as a consequence of eq.(55) the variation of the temperature and luminosity of the black hole depends on the sign of $\partial_{t}\Lambda$ .
+If eqs.(48)-(49) are applied in eq.(25) then
+$\begin{array}[]{l}\frac{3}{2}M\partial_{t}G+G\partial_{t}M+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (57)
+If $\partial_{t}\Lambda=0$ then from eq.(57) we obtain $\gamma=\frac{2}{3}$.
+Surprisingly, this is the same relation between the Newton constant and the mass which follows from the formula for the mass of the neutron star and its Chandrasekhar limit [45] (sec.11) According to ref.[18] it could serve as an estimate of the variation of the Newton constant.
+Let us consider the general option (43) with $\gamma\neq 0$ . When we substitute the relation (43) between $G$ and $M$ in eq.(53) then we obtain the solution for $M$ ( when $\gamma\neq\frac{3}{2}$)
+$M=M_{0}\Big(1-(3-2\gamma)\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{1}{3-2\gamma}}\equiv M_{0}Z^{\frac{1}{3-2\gamma}},$ (58)
+$G=G_{0}\Big(1-(3-2\gamma)\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{\gamma}{3-2\gamma}}\equiv G_{0}Z^{-\frac{\gamma}{3-2\gamma}}.$ (59)
+When $\gamma=\frac{3}{2}$ then
+$M=M_{0}\exp\Big(-\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)$ (60)
+It follows that ( for $\gamma\neq\frac{3}{2}$)
+$\begin{array}[]{l}GM=G_{0}M_{0}Z^{\frac{1-\gamma}{3-2\gamma}}\end{array}$ (61)
+$\begin{array}[]{l}GM^{2}=G_{0}M_{0}^{2}Z^{\frac{2-\gamma}{3-2\gamma}}\end{array}$ (62)
+and
+$\begin{array}[]{l}G^{-3}M^{-2}=G_{0}^{-3}M_{0}^{-2}Z^{\frac{3\gamma-2}{3-2\gamma}}\end{array}$ (63)
+These equations have some consequences for the behaviour of the temperature, luminosity (L), entropy and the mean density (defined as the mass divided by the volume).
+Clearly, $M$ tends to zero for all $\gamma$ and $G$ tends to infinity.
+However, the dependence of the temperature, luminosity, entropy and the mean desity may substantially depend on $\gamma$.
+For the temperature and luminosity we have
+$\gamma$ T,L time
+$0<\gamma<1$ increasing to infinity finite
+$\gamma=1$ constant all time
+$1<\gamma<\frac{3}{2}$ decreasing to zero finite
+$\gamma\geq\frac{3}{2}$ decreasing to zero infinite
+For $\gamma\geq\frac{3}{2}$ the mass, the temperature and the luminosity tend to zero asymptotically in time.
+Let us note that the entropy (62) is decreasing to zero if $\gamma<2$, is constant if $\gamma=2$ and is growing in time if $\gamma>2$.
+The mean density (63)is tending to infinity if $\gamma<\frac{2}{3}$, is constant for $\gamma=\frac{2}{3}$ and is tending to zero if $\gamma>\frac{2}{3}$ (this is different than in the Hawking-Page model [72]).
+The results could be compared with observations of the gamma-ray bursts [26].
 In particular, with the recent observations of the giant neutrino emissions [25] which by some authors [23][24] are interpreted as resulting from primordial black hole explosions.
-The life time of the black hole according to eq.(43) is
-$t_{f}=\frac{3}{5}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (47)
-It is longer by a factor $\frac{9}{5}$ than the one of the Hawking-Page model (which would result from the assumption that $G=const$ in eq.(39)).
-5 Summary and outlook
-We have shown that the Bianchi identity of classical Einstein equations admits a violation of the ideal fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time.
-As a generalization of the first and the second laws of thermodynamics the covariant energy-momentum divergence can be expressed by the covariant divergence of the entropy density.
-Then, the entropy density enters the Bianchi identity.
-In this paper we have applied the formula for the change of entropy during the black hole evaporation to determine the time-dependence of $G$ in the Bianchi identity.
-As a consequence of a growing $G$ the black hole evaporation is slower and the average black hole density is constant in contradistinction to the one for a fixed $G$ (this could be subject to astronomical observations) .
-The time-dependence of $G$ is unavoidable if we wish to describe the non-conserved energy-momentum of an evaporating black hole by Einstein equations (2).
-The increasing $G$ could have a physical stabilizing origin preventing the appearance of an infinite density when the black hole’s radius tends to zero.
-Another problem deserving a study concerns the entropy of an expanding universe with a cosmological horizon.
-For such a system the varying cosmological constant in the Bianchi identity should be relevant.
+The life time of the black hole for $\gamma<\frac{3}{2}$ according to eq.(58) is
+$t_{f}=\frac{1}{3-2\gamma}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (64)
+It is infinite for $\gamma\geq\frac{3}{2}$ but the mass is decreasing asymptotically to zero.
+6 Summary
+We have shown that the Bianchi identity of classical Einstein equations admits a violation of the fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time.
+The covariant energy-momentum divergence of the fluid appearing in the Bianchi identity has the form of the heat flow.
+We define the mass of the fluid as a volume integral of its energy density.
+If the pressure of the system is zero then the Bianchi identity relates $\partial_{t}(GM)$ to the time derivative of the cosmological term.
+For a non-zero pressure an analogous formula involves the temperature and entropy.
+A radiation of the fluid (involving an energy loss) can be included in Einstein equations if the the Newton “constant” or the cosmological term depend on time.
+The radiation is described by the Stefan-Boltzmann law.
+The Stefan-Boltzmann formula depends on the temperature and the area of the emitting surface.
+When the collapse of the star is described by pasting of Schwarzschild solution with the FLWR solution for a fluid with a certain density, then the resulting black hole should have the mass $M$ related to this density.
+We use the mass $M$ for a determination of the temperature and the area.
+The black hole mechanics derived in classical Einstein gravity gives a relation between $G$, $M$, the entropy $S$ and temperature $T$.
+In some effective theories of gravity when $G$ is time-dependent this relation may be more involved.
+From the Bianchi identity we obtain an equation relating the time derivatives of $G$, $\Lambda$ and $M$ to the temperature $T$ and entropy $S$.
+We discussed various formulas for $T$ and $S$ some of them directly generalizing the Bekenstein-Hawking expression and some (for $\gamma\neq 1$ and $\gamma\neq\frac{2}{3}$) quite hypothetical.
+These assumptions give a relation $G\simeq M^{-\gamma}$.
+When the expression $G(M)$ is substituted in the Stefan-Boltzmann formula for the mass decrease then it gives an evaporation law without free parameters.
+Depending on the value of $\gamma$ the black hole temperature and luminosity may either tend to infinity at finite time (black hole explosion), be constant during the whole black hole life time or even tend to zero in finite or infinite time.
+The latter option is interesting for the proposals of the dark matter as consisting of primordial black holes.
+The study of the possible black holes evaporation laws should give useful hints for their search in astronomical observations.
 Data availability statement No data were created or analysed in this study
-Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis. JH thanks Dr. Aleksander Kozak for supervising the thesis.
+Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis.
+JH thanks Dr.
+Aleksander Kozak for supervising the thesis.
+ZH thanks Hytham Alwrekat for discussions on some topics of this paper.